OpenStudy (gabylovesyou):

Using the following equation, find the center and radius of the circle. You must show all work and calculations to receive credit. x2 + 2x + y2 + 4y = 20 @amistre64

OpenStudy (amistre64):

i beleive you want to do thru a process called completing the square

OpenStudy (amistre64):

you will also want to recall what the general equation of a circle looks like

OpenStudy (amistre64):

yep, thats the general equation, and completing the square with that we have will transform it into that format

OpenStudy (gabylovesyou):

ok how do we do that

OpenStudy (amistre64):

well it not that hard to do: a complete square looks like:\[x^2 + bx + b^2/4\]

OpenStudy (amistre64):

the reason it looks that way tends to be more useful than why it looks that way

OpenStudy (amistre64):

(x+n)^2 = (x+n)(x+n) = x^2 +2n x + n^2 as such: n^2 = (2n/2)^2

OpenStudy (amistre64):

since we want the (x+n) part of it, then when 2n = b n=b/2

OpenStudy (gabylovesyou):

hmmmm ;/ im trying to understand this

OpenStudy (amistre64):

lots of practice helps :)

OpenStudy (amistre64):

x2 + 2x + y2 + 4y = 20 (x2 + 2x) + (y2 + 4y) = 20 consider: x^2 + 2n x + n^2 = ax^2 + b x + c 2n = b, n = b/2, n^2 = b^2/4

OpenStudy (amistre64):

yes, and so far i have attempted to describe up to step 3

OpenStudy (amistre64):

complete the square is what i have always known it as

OpenStudy (amistre64):

(x2 + 2x) + (y2 + 4y) = 20 +1 +4 +1+4 ---------------------------- (x2 +2x+1) + (y2 +4y+4) = 25

OpenStudy (amistre64):

(2/2)^2 = 1 and (4/2)^2 = 4 as i made mention of at the start

OpenStudy (amistre64):

but the key is in knowning how to condense the sqaures into an (x+n)^2 format n = b/2; so nx = 2/2 and ny = 4/2

OpenStudy (gabylovesyou):

ok..

OpenStudy (gabylovesyou):

i dont think ima learn how to do that now... so how do i find the radius and center ?

OpenStudy (amistre64):

you have to do that in order to find them. if you cant learn that at the moment then you will have a rather difficult time finding them

OpenStudy (amistre64):

one simple method would be to use the wolf, itll give you radius and center in its properties list

OpenStudy (amistre64):

(x + 1)^2 + (y + 2)^2 = 25 ; yes

OpenStudy (amistre64):

r^2 = 25 so r = sqrt(25) and you already stated you h and k parts effectively to me

OpenStudy (gabylovesyou):

so good ? haha

OpenStudy (gabylovesyou):

or is it 5 ..

OpenStudy (amistre64):

yeah, its good :) just work the steps they gave you with the values required r = sqrt(25) = 5

OpenStudy (gabylovesyou):

thats the radius ?

OpenStudy (amistre64):

i hope so, otherwise im trying to figur eout what temperature to set the dishwasher at

OpenStudy (gabylovesyou):

cx lol ok thanks.

OpenStudy (amistre64):

good luck :)